UARS Weekly Status Report

17 May 2002


                UARS WEEKLY STATUS REPORT - 17 MAY 2002



SOLSTICE continues under day only operations due to low power
conditions. We are performing default solar observations during this


SUSIM successfully performed its daily mid-resolution and MgII
solar scans during periods of predicted UARS telemetry.

We have made adjustments to our solar pointing correction
model and expect to reprocess the entire data set soon.


Nothing received.


PEM wishes to thank the SUSIM team for producing UARS attitude files.
The PEM team was able to up-load these data to CDHFB:: and generate the PEM
O/A subset information which is allowing PEM data from greater than 30 keV 
electrons and protons to be correctly generated.

PEM support of the UARS Traceability Mission for the TIMED mission has
been organized with the GUVI team.  PEM data was collected from April 18
through April 24.  During this time period, there were two spacecraft contacts. 
Early PEM data from the contact times showed no aurora.  No significant aurora
appeared at the predicted highest invariant latitudes; however, lower invariant
latitude passes did show significant aurora.  It appears as though magnetic
activity increased at times when UARS was not at its peak invariant latitudes
and some of the particle precipitation activity was captured by the PEM 
instrumentation.  PEM wishes to thank Bob Neff and Rich Beck at GSFC for their 
efforts in retriving as much of the PEM data as possible.

Currently, PEM specific orbit/attitude and high energy particle data is 
being processed for the GUVI-PEM Traceability study.  Evaluation of PEM data 
and spacecraft locations will need to be assimilated.  These data are generated
so that the PEM data may be mapped on to GUVI images.  Currently, a selected 
set of PEM images for electrons less than 30 keV showing auroral activity may
be found at "".  Magnetometer DC Bz
and AX Bx are shown to aid in identifying the point of highest invariant 
latitude.  The solar array current and voltage are shown to aid in determining 
when the spacecraft is in sunlight.

PEM has provided input to the "Operations Agreement (OA) Between the 
Packet Processor-Automation (Pacor-A) and Particle Environment Monitor (PEM)
for Upper Atmosphere Research Satellite (UARS)".  This document is part of the 
Consolidated Space Operations Contract.

PEM continues to receive NMC data files from NOAA.  No problems have
been encountered.

*Beginning of power orbit day (PEM activation): array current >5 amps, array 
     voltage >53.5 volts, and battery current >2.0 amps.  
 End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
     array current <9 amps, and battery currents <0.2 amps.
 Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.


HALOE continues to operate successfully after powering on May 11, 
2002. All engineering parameters are within nominal operating 
region. Some problems, however,  were encountered powering 
HALOE on that were exacerbated by lack of a real-time KCRT.

With use of the new ECLIPSE system there is no provision for real-time 
instrument KCRT connections. This creates an extra burden on both the 
instrument FOT and the UARS FOD especially during power on and special 
operations. The FOD has to monitor the HALOE pages for us and verbally 
provide parameter information. This process did not work well this past 
power on for HALOE. The new eclipse machine is much faster than the old 
AP's that were formally used and a bug in the code that worked before 
caused ground commands and RTS's to be executed and sent to HALOE at the 
same time. Several attempts to power on HALOE on Friday May 10 were 
unsuccessful. Finally, on the following day with Dimitrios supporting the 
FOD by closely monitoring the power on process he found the operational bug 
and a subsequent power on was successful. If we had a KCRT at LaRC the 
problem could possibly have been diagnosed much sooner. We do appreciate, 
however, the diligent efforts by Dimitrios and Jim Elliot in getting HALOE 
back on.

Additionally, we were initially delayed powering-on by a special 
battery-charging test conducted by the UARS FOT. As much as HALOE is off, 
this test could have been conducted on a non-interfering basis. Also the 
use of the tape recorder was suspended to support the test resulting in the 
potential loss of six out of the eight possible events for the day if HALOE 
had successfully turned on initially.


Nothing received.


Nothing received.


Nothing received.


The MPG participated in a meeting with the UARS and TIMED
Project Scientists on 15 May to coordinate operations scheduling
of UARS instruments.

The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.

Daily processing of instrument Daily Activity Plans continues
with no problems.


Nothing received.

                          REPORTING PERIOD

This report covers 27 April 2002 (Orbit 58100, GMT Day 117) through 03
May 2002 (Orbit 58204, GMT Day 123).

                         SPACECRAFT OPERATION

The observatory is now in Forward Flight (Southern Hemisphere viewing)
and   is  performing  nominally  under  two-battery  operations.   The
instrument operational changes were:


       Instrument    Time          Orbit     Comment
       HALOE         117/0237      58101     HALOE POWER OFF
       HRDI          122/1432      58184     HRDI TO LOW POWER MODE

The  beta  angle  ranged between the angles of 53.74 degrees  to  72.6
degrees  for  this  report period.  The beta angle is  now  decreasing
towards a minimum angle of 0 degrees on 23 May 2002 (DOY 143).

Spacecraft  battery  1,  2,  and 3 performance  monitoring  continues.
Battery 1 remains OFF the charge relay.  Battery 2 and Battery  3  are
currently maintaining greater than 23.8 V end-of-night (EON) load  bus
voltage.   Battery  temperatures are stable with a  temperature  delta
between Battery 2 and Battery 3 of 0.95 to 0.96 degrees C.  Additional
battery performance data for each battery on 26 April and 03 May  2002
are provided in Appendix D.

The  clock  error  ranged from -10.3 to 11.5 msec during  this  report
period.  Clock rate adjustments are listed in Appendix C.

All  instrument  and subsystem engineers were advised on  the  initial
excursion  of ALERTS or Out-Of-Limits conditions detected by  the  FOT
for  this report.  A summary is listed below and the details of  ALERT
and Out-Of-Limits occurrences are listed in Appendix B.

    OUT-OF-LIMITS  =    00
    ALERTS         =    00

The  following real-time command sheets were executed for  the  listed
operational element(s) during this report period (see Appendix A).

    FOT            =    05
    SUSIM          =    02
    WINDII         =    01
    HALOE          =    01
    HRDI           =    01

UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.

    FOT       =    4

Four  (4)  Anomaly  Reports remained OPEN at the end  of  this  report
period (see Appendix E).

    Transponder A:  01/333  - 2287.496855
    Transponder B:  02/043  - 2287.498879

                        GROUND SYSTEM OPERATION

Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.

The data loss  calculations are no longer being supplied by Data
Capture Facility (DCF).

                       OTHER SIGNIFICANT EVENTS

                       UARS SATELLITE OPERATIONS

The  FOT is in the process of implementing some minor changes  to  the
Eclipse  system  to  help optimize operations.  All  of  the  existing
documentation is also in the process of being updated to  reflect  the
new  ground  system  nomenclature.  A few new  SOPs  have  yet  to  be
written, while the handful that have already been written need  to  be
finalized and signed off on.

A  discussion  regarding the UARS power issues was held  at  the  UARS
Monthly meeting on 02 May 2002.  Further discussion and data is needed
before  a decision will be made.  A follow up meeting will be held  on
09 May 2002.

TMON 1 tripped on 03 May 2002 (123) at approximately 06:40 GMT as  the
LBV  dropped below 23.6 due to high beta.  As a result of  this  HRDI,
PEM and WINDII were powered off.  PEM powered up on its own because it
was  in Day-Only mode.  HRDI was powered back on and in Day-Only  mode
on 126/1505.  WINDII resumed operations on 127/1826.


The FOT attended the UARS Monthly Meeting on 02 May 2002.

                       FUTURE SIGNIFICANT EVENTS

The  next UARS Yaw slew is scheduled for 23 May 2002 (DOY 143).   This
maneuver  will  be  a  Forward to Reverse  Flight  maneuver  (Northern
Hemisphere viewing).

Attachments:      Appendix A,              Real-time Command Sheets
                  Appendix B,              Out-Of-Limits   &   ALERT
Occurrences       Appendix C,              Clock &  Solar  Array  Rate

Corrections       Appendix D,              UARS Battery Performance Data
                  Appendix E,              UARS Anomaly Reports
                            APPENDIX A
                       REAL-TIME COMMAND SHEETS
                        Week ending 03 May 2002

ORBIT   AOS        NO.      SUBSYS      REASON
-----   ---        ---      ------      ------
58101   117/0237   02-148   HALOE       POWER OFF
58113   117/2203   02-149   OBC/CLK     CLK ADJ FR 20.6 TO 0.0
58134   119/0640   02-150   SUSIM       SUSIM CMDS
58134   119/0738   02-152   OBC/CLK     CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
58135   119/0918   02-151   SUSIM       SUSIM CMDS
58138   119/1248   02-151   SUSIM       SUSIM CMDS
58139   119/1435   02-147   WINDII      WINDII RESET ERR E06
58145   120/0048   02-147   WINDII      WINDII RESET ERR E06
58148   120/0556   02-153   OBC/CLK     CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
58160   121/0048   02-147   WINDII      WINDII RESET ERR E06
58175   122/0048   02-147   WINDII      WINDII RESET ERR E06
58178   122/0553   02-154   OBC/CLK     CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
58184   122/1432   02-155   HRDI        POWER LOW POWER MODE
58187   122/1932   02-156   OBC         STATBUF DUMP
                              APPENDIX B
                        Week ending 03 May 2002

The following Out-Of-Limits occurrences were detected:

--------       -----   -----------                ------
-None this report period.

The following ALERT occurrences were detected:

MNEMONIC     DESCRIPTION                          ORBITS
--------     -----------                          ------
-None this report period.


                              APPENDIX C
                        Week ending 03 May 2002


ORBIT          TIME                   CLOCK RATE CHANGE
-----          -----                  -----------------
58113          117/2205               CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
58134          119/0738               CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
58148          120/0602               CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
58178          122/0612               CLK ADJ FR 41.2 TO 20.6 MSEC/DAY

                              APPENDIX D
                        Week ending 03 May 2002

     26 April 2002 - GMT Day 116 Beta = 53.74 deg, SA Pos = 269 deg

      DIFF V (mV)  TEMP (C)     CURR (amp)   EON V    AVE C/D  AVE
BAT   MAX/MIN      MAX/MIN      MAX/MIN      MIN      RATIO    DOD(%)
---   -------      --------     ----------   -----    -------  ------
1   +728.0/+728.0  -3.16/-4.38  +0.0/ +0.0   0.0      0.0      0.0
2   +11.2/-100.8   +5.14/+3.87  +24.0/-15.2  24.8     1.040   20.2
3   +22.4/-56.0    +3.87/+2.91  +24.0/-15.2  24.8     1.040    19.9

     03 May 2002 - GMT Day 123 Beta = 72.6 deg, SA Pos = 269 deg

     DIFF V (mV) TEMP (C)     CURR (amp)  EON V    AVE C/D  AVE
BAT  MAX/MIN     MAX/MIN      MAX/MIN     MIN      RATIO    DOD(%)
---  -------     --------     ----------  -----    -------  ------
1  +728.0/+728.0             -3.16/-4.38  +0.0/ +0.0         0.0 0.000
2   +11.2/-44.8 +2.60/+1.65  +16.8/-14.8  23.5     1.110    13.7
3   +15.5/-44.8 +1.97/+1.02  +16.0/-14.8  23.4     1.110    13.5

ORB#           TIME       LVL FR-TO
-----          ----       ---------
-no changes    V/T 5

                              APPENDIX E
                         UARS ANOMALY REPORTS
                        Week ending 03 May 2002

------    ------    -----     ----      ---       -----------------
02-026    FOT       58172     121/2016  14170     Generic Late Acquisition
02-026    FOT       58192     123/0307  14170     Generic Late Acquisition
02-026    FOT       58195     123/0954  14170     Generic Late Acquisition
02-027    FOT       58202     123/1931  24324     Late Acquisition

------    ------    -----     ----      ---       -----------------
02-025    FOT       See List  Apr 13    N/A       Generic Late Acquisition

------    ------    -----     ----      ---       -----------------
-None this report period.

                         OPEN ANOMALY REPORTS
------    ------    -----     ----      ---       -----------------
02-004    SOLSTICE  56883     036/0050  N/A       Solstice in Wrong Mode
02-023    FOT       57767     094/2027  24285     Early LOS on OMNI Pass
02-024    FOT       57783     095/2209  24285     SOLSTICE INOPS in Day
02-027    FOT       58202     123/1931  24324     Late Acquisition